# Sentiment Without Structure Differential Market Responses to Infrastructure vs Regulatory Events in Cryptocurrency Markets

### v1

@@ Abstract
We investigate differential market responses to infrastructure versus regulatory events in cryptocurrency markets using event study methodology with 4-category event classification. From 50 candidate events (2019–2025), 31 meet inclusion criteria across Bitcoin, Ethereum, Solana, and Cardano. We employ constant mean and market-adjusted models with event-level block bootstrap confidence intervals that properly account for cross-sectional correlation.
Our primary comparison focuses on negative-valence events: infrastructure failures (8 events) versus regulatory enforcement (7 events). Infrastructure failures produce mean Cumulative Abnormal Return (CAR) of $7.6\%$ (95% CI: $25.8\%,+11.3\%]$) and regulatory enforcement produces mean CAR of $11.1\%$ (CI: $31.0\%,+10.7\%]$). The difference of +3.6+3.6 percentage points has CI $25.3\%,+30.9\%]$, p=0.81p=0.81—a null finding indicating markets respond similarly to both shock types when controlling for valence.
Robustness checks confirm consistent results across window specifications, leave-one-out exclusion of major events (FTX, Terra), and alternative market model specifications. The 4-category classification addresses prior conflation of upgrades with failures. This exploratory analysis should be treated as hypothesis-generating.
Keywords: Cryptocurrency, Event Study, Regulation, Infrastructure Risk, Block Bootstrap
JEL Codes: G14, G18, G23
@@ 1 Introduction
Do cryptocurrency markets respond differently to infrastructure failures versus regulatory enforcement? This paper investigates differential market responses using a 4-category event classification that addresses prior conflation of positive and negative events within each type. This study extends prior work on asymmetric volatility response to cryptocurrency shocks (Farzulla, 2025) by focusing on return-level comparisons rather than conditional variance dynamics.
@@ 2.1 Event Studies in Cryptocurrency Markets
Event study methodology has been applied to cryptocurrency markets by Auer and Claessens (2018) and Lucey et al. (2022), with comprehensive surveys provided by Kothari and Warner (2007). The foundational event study framework established by MacKinlay (1997) has been adapted to cryptocurrency settings, though these studies typically find significant abnormal returns around major announcements but rarely distinguish between event types systematically. Recent work has examined regulatory impacts across multiple jurisdictions (Chokor and Alfieri, 2021; Koenraadt and Leung, 2024; Cumming et al., 2025), generally finding negative market reactions to restrictive regulation.
@@ 2.2 Market Microstructure
Cryptocurrency microstructure differs from traditional markets (Makarov and Schoar, 2020). Aste (2019) provides foundational analysis of dependency and causality structures across nearly two thousand cryptocurrencies, demonstrating that prices are significantly correlated with social sentiment and that major assets like Bitcoin occupy central positions in price correlation networks while playing marginal roles in sentiment networks. This asymmetry between price and sentiment network centrality motivates our focus on how different shock types propagate through market structure. The development of systematic benchmarks such as CRIX (Trimborn and Härdle, 2018) has enabled rigorous market-wide analysis, while indices capturing regulatory-specific risk such as CRRIX (Ni et al., 2025) demonstrate that policy news systematically impacts crypto market dynamics. The absence of designated market makers, 24/7 trading, and cross-exchange fragmentation complicate event study inference. Exchange hacks and regulatory news have been shown to significantly impact Bitcoin volatility (Lyócsa et al., 2020), with spillover effects across international markets (Borri and Shakhnov, 2020). We focus on returns, leaving liquidity extensions to future work.
@@ 2.3 Infrastructure vs Regulatory Distinction
Infrastructure events: Direct mechanical impact on transaction execution, custody, protocol functionality, or exchange operations. Examples: exchange failures (Mt. Gox, FTX) with documented systemic contagion effects (Jalan and Matkovskyy, 2023; Conlon et al., 2024), protocol failures such as the Terra/UST algorithmic stablecoin collapse (Liu et al., 2023; Badev and Watsky, 2023), and network incidents.
@@ 3.3 Price Data
Daily open-high-low-close-volume (OHLCV) data from Binance (2019-01-01 to 2026-01-29). Returns computed as log-differences of closing prices.
@@ 4.2 Return Models
Capped Returns. Following standard practice for cryptocurrency event studies, we winsorize daily returns at $\pm 50\%$ to limit the influence of extreme outliers while preserving economically meaningful variation. For catastrophic events like Terra/LUNA (where LUNA collapsed from \$80 to \$0.01), uncapped returns would produce CARs exceeding $1000\%$, violating standard regression assumptions and dominating all statistical inference. Capping at $\pm 50\%$ per day retains these events in the sample while maintaining statistical validity:
@@
$AR_{it}=\tilde{R}_{it}-\bar{R}_{i}$ (2)
where $AR_{it}$ denotes the Abnormal Return for asset $i$ on day $t$, and $\bar{R}_{i}$ is estimated from the 250-day estimation window using capped returns.
@@ 5.11 Winsorization Sensitivity
To address concerns about the $\pm 50\%$ return cap, we re-estimate results across four winsorization levels (Table 10).
@@ 6.1 What the Data Actually Show
Interpretation: We cannot distinguish between the two event types statistically. Both produce similar negative returns. The enforcement capacity hypothesis is not confirmed, but neither is the alternative that regulatory events have larger effects.
@@ 6.2 Theoretical Framework: Bounded vs Unbounded Uncertainty
  • Timeline bounded: Infrastructure crises typically resolve within weeks to months.
@@
  • Timeline unbounded: Regulatory uncertainty can persist indefinitely—the SEC has never provided clear classification criteria.
Implication for similar CARs: If markets price both shock types similarly despite different mechanisms, it may reflect: (1) offsetting effects where bounded infrastructure losses are acute but recoverable, while unbounded regulatory uncertainty persists but is less immediately devastating, or (2) market participants simply responding to negative sentiment regardless of source.
@@ 6.4 Implications
For investors: Neither event type clearly dominates in market impact when controlling for valence. Both infrastructure failures and regulatory enforcement produce roughly 8–11% negative CARs.
@@ 7 Conclusion
What this means: The enforcement capacity hypothesis—that infrastructure events should dominate regulatory events—is neither confirmed nor rejected. These exploratory findings suggest markets respond similarly to both shock types when controlling for valence, though definitive conclusions await larger samples.
@@
Future research: Larger event samples (power analysis suggests $N\approx 3{,}000$ per category for 80% power), pre-registered hypotheses, non-overlapping event windows, and multi-venue price data are needed for confirmatory tests. The “uncertainty resolution” interpretation (Section 6.2) remains speculative pending prospective testing. Named-token analysis à la Saggu et al. (2025)—examining directly treated assets (e.g., XRP in SEC v. Ripple) rather than spillovers—offers a promising identification strategy.

### v2

# Dissensus AI Working Paper Series DAI-2507

@@ Abstract
We investigate whether cryptocurrency markets differentiate between infrastructure failures and regulatory enforcement at the return level, complementing a companion conditional variance analysis that finds 5.7 times larger volatility impacts from infrastructure events ($p=0.0008$). Using event-level block bootstrap inference on 31 events across Bitcoin, Ethereum, Solana, and Cardano (2019–2025), we find no statistically significant difference in cumulative abnormal returns between infrastructure failures ($7.6\%$) and regulatory enforcement ($11.1\%$): the difference of +3.6+3.6 pp has $p=0.81$ with 95% CI $25.3\%,+30.9\%]$.
This null acquires substantive meaning alongside the companion’s highly significant variance result: the same events that produce indistinguishable return responses generate dramatically different volatility signatures. Markets differentiate shock types through the risk channel—the second moment—rather than expected returns. The block bootstrap methodology, which resamples entire events to preserve cross-sectional correlation, reveals that prior parametric approaches systematically understate uncertainty by inflating degrees of freedom. Results are robust across eight specifications including permutation tests, leave-one-out analysis, and the Ibragimov–Müller few-cluster test.
Keywords: Cryptocurrency, Event Study, Block Bootstrap, Volatility Asymmetry, Infrastructure Risk, Multi-Moment Analysis
JEL Classification: C22, C58, G12, G14
@@ 1 Introduction
Do cryptocurrency markets respond differently to infrastructure failures versus regulatory enforcement? This paper investigates differential market responses using a 4-category event classification that addresses prior conflation of positive and negative events within each type. This study extends prior work on asymmetric volatility response to cryptocurrency shocks (Farzulla, 2025a) by focusing on return-level comparisons rather than conditional variance dynamics. That companion study finds infrastructure events generate 5.7 times larger conditional variance increases than regulatory events ($p=0.0008$, Cohen’s $d=2.753$) using GJR-GARCH-X models. A natural follow-up question—addressed here—is whether this asymmetry extends to the first moment: do infrastructure failures also produce larger returns than regulatory enforcement, or does the differential response operate exclusively through the risk channel?
@@ 2.1 Event Studies in Cryptocurrency Markets
Event study methodology has been applied to cryptocurrency markets by Auer and Claessens (2018) and Lucey et al. (2022), with comprehensive surveys provided by Kothari and Warner (2007) and systematic reviews of social media sentiment influence by Sundarasen and Saleem (2025). The foundational event study framework established by MacKinlay (1997) has been adapted to cryptocurrency settings, though these studies typically find significant abnormal returns around major announcements but rarely distinguish between event types systematically. Recent work has examined regulatory impacts across multiple jurisdictions (Chokor and Alfieri, 2021; Koenraadt and Leung, 2024; Cumming et al., 2025), generally finding negative market reactions to restrictive regulation.
@@ 2.2 Market Microstructure
Cryptocurrency microstructure differs from traditional markets (Makarov and Schoar, 2020). Aste (2019) provides foundational analysis of dependency and causality structures across nearly two thousand cryptocurrencies, demonstrating that prices are significantly correlated with social sentiment and that major assets like Bitcoin occupy central positions in price correlation networks while playing marginal roles in sentiment networks. This asymmetry between price and sentiment network centrality motivates our focus on how different shock types propagate through market structure. Transfer entropy analysis further reveals that BTC serves as the dominant information driver in crypto-equity interactions, with announcement-driven feedback from stocks to BTC occurring only during major financial events (Aufiero et al., 2025a). The development of systematic benchmarks such as CRIX (Trimborn and Härdle, 2018) has enabled rigorous market-wide analysis, while indices capturing regulatory-specific risk such as CRRIX (Ni et al., 2025) demonstrate that policy news systematically impacts crypto market dynamics. The absence of designated market makers, 24/7 trading, and cross-exchange fragmentation complicate event study inference. Exchange hacks and regulatory news have been shown to significantly impact Bitcoin volatility (Lyócsa et al., 2020), with spillover effects across international markets (Borri and Shakhnov, 2020). Sentiment effects extend beyond spot returns: Kim (2025) finds that social media peer opinions significantly predict cryptocurrency option prices, suggesting that the information channel operates across multiple market layers. Meanwhile, Ateş and Başarslan (2025) compare traditional and contextual NLP representations for cryptocurrency sentiment, finding inconsistent performance across methods—results consistent with our premise that structural assumptions may impose false patterns rather than capturing genuine signal. We focus on returns, leaving liquidity extensions to future work.
@@ 2.3 Infrastructure vs Regulatory Distinction
Infrastructure events: Direct mechanical impact on transaction execution, custody, protocol functionality, or exchange operations. Examples: exchange failures (Mt. Gox, FTX) with documented systemic contagion effects (Jalan and Matkovskyy, 2023; Conlon et al., 2024), protocol failures such as the Terra/UST algorithmic stablecoin collapse (Liu et al., 2023; Badev and Watsky, 2023), and network incidents. Developer-level evidence corroborates this distinction: Vaccargiu et al. (2025) find that technical events in Ethereum repositories generate anticipatory development activity before events followed by reduced commit rates, whereas market events produce reactive development patterns—mirroring the anticipation-reaction asymmetry we document in price dynamics.
@@ 3.3 Price Data
Daily open-high-low-close-volume (OHLCV) data from Binance (2019-01-01 to 2026-01-29). Returns computed as percentage changes of closing prices.
@@ 4.2 Return Models
Return Capping. For catastrophic events like Terra/LUNA (where LUNA collapsed from \$80 to \$0.01), uncapped returns would produce CARs exceeding $1000\%$, violating standard regression assumptions and dominating all statistical inference. Following standard practice for cryptocurrency event studies, we define capped returns as:
@@
Our primary analysis uses uncapped returns ($R_{it}$). Winsorization at $\pm 50\%$ and other cap levels is applied as a robustness check (Section 5.11), where results are virtually identical across all specifications (see Table 10).
@@
$AR_{it}=R_{it}-\bar{R}_{i}$ (2)
where $AR_{it}$ denotes the Abnormal Return for asset $i$ on day $t$, and $\bar{R}_{i}$ is estimated from the 250-day estimation window.
@@ 5.11 Winsorization Sensitivity
To assess sensitivity to extreme returns, we re-estimate results across four winsorization levels (Table 10).
@@ 6.1 What the Data Actually Show
[…]
@@ 6.2 Theoretical Framework: Bounded vs Unbounded Uncertainty
  • Timeline bounded: Infrastructure crises typically resolve within weeks to months. The companion volatility study confirms this pattern empirically: infrastructure sensitivity amplifies fivefold during crisis regimes but reverts as crises resolve, while regulatory effects remain temporally flat (Farzulla, 2025a).
@@
  • Timeline unbounded: Regulatory uncertainty can persist indefinitely—the SEC has never provided clear classification criteria. Consistent with this unbounded character, the companion study finds regulatory volatility coefficients are stable across calm and crisis regimes (0.51% vs 0.42%)—regulatory uncertainty operates as a persistent baseline rather than an acute spike (Farzulla, 2025a).
Implication for similar CARs: If markets price both shock types similarly despite different mechanisms, it may reflect: (1) offsetting effects where bounded infrastructure losses are acute but recoverable, while unbounded regulatory uncertainty persists but is less immediately devastating, or (2) market participants simply responding to negative sentiment regardless of source. The latter interpretation is consistent with evidence that the information content of cryptocurrency project fundamentals has limited predictive power for market behavior (Farzulla, 2025b), suggesting that sentiment-driven pricing may dominate structural analysis across multiple information channels.
@@
6.3 Companion Evidence: The Structure Lives in the Second Moment
The null finding at the return level acquires interpretive weight when read alongside the companion study’s highly significant variance result (Farzulla, 2025a). Infrastructure events produce similar mean CARs to regulatory events ($7.6\%$ vs $11.1\%$, $p=0.81$) but generate 5.7 times larger conditional variance increases ($2.385\%$ vs $0.419\%$, $p=0.0008$). This pattern implies that the market’s differential response to infrastructure versus regulatory shocks operates through risk—the second moment—rather than expected returns.
This decomposition is economically coherent. Both shock types carry negative sentiment and produce directionally similar price declines. However, infrastructure failures create immediate mechanical disruptions (exchange halts, protocol exploits, liquidity cascades) that generate sharp variance spikes, while regulatory events operate through information channels requiring gradual interpretation. The companion study further shows infrastructure sensitivity amplifies fivefold during crisis regimes ($F=45.23$, $p<0.001$), while regulatory effects remain flat across regimes—a non-linearity invisible to return-level analysis.
The Knight framework developed above is consistent with this variance-domain evidence: infrastructure events represent bounded, quantifiable shocks that generate acute but time-limited variance spikes, while regulatory events introduce persistent but lower-magnitude uncertainty. Together, the two papers suggest that traditional event studies focusing solely on CARs systematically miss the risk channel through which markets differentially process shock types.
Methodological note on companion differences: The companion study uses CoinGecko-sourced prices for six cryptocurrencies (BTC, ETH, XRP, BNB, LTC, ADA) over January 2019–August 2025 with log returns, while the present study uses Binance-sourced prices for four cryptocurrencies (BTC, ETH, SOL, ADA) over January 2019–January 2026 with percentage returns. The overlapping subset (BTC, ETH, ADA; January 2019–August 2025) enables cross-study comparison, though differences in data source, return definition, and asset universe should be noted. The companion study’s binary event classification (infrastructure vs regulatory) maps directly onto our 4-category framework’s negative-valence subset.
@@ 6.5 Implications
[…]
@@ 7 Conclusion
What this means: The enforcement capacity hypothesis—that infrastructure events should dominate regulatory events—is neither confirmed nor rejected at the return level. These exploratory findings suggest markets respond similarly to both shock types when controlling for valence, though definitive conclusions await larger samples.
[…]
@@
Future research: Larger event samples (power analysis suggests $N\approx 930$ per category for 80% power at the observed effect size), pre-registered hypotheses, non-overlapping event windows, and multi-venue price data are needed for confirmatory tests. The “uncertainty resolution” interpretation (Section 6.2) remains speculative pending prospective testing. Named-token analysis à la Saggu et al. (2025)—examining directly treated assets (e.g., XRP in SEC v. Ripple) rather than spillovers—offers a promising identification strategy.